A 9.0 V battery and two restitors are connected. The current through the 3.0 ohm resistor is 2.0 A. What is the potential difference across the restiror R2
1
Expert's answer
2018-10-05T14:35:09-0400
Obviously, the condition implies a series connection of resistors
Resistance:
R_1=3Ohm
Voltage:
U=9V
Current:
I=2A
Solution:
According to Ohm's law, the voltage drop across the first resistor:
U_1=R_1 I=3*2=6V
U_2=U-U_1=9-6=3V
Answer:
The voltage drop across the R2 resistor is:
U_2=3V